Find the future value of the following investment.

Principal

Nominal Rate

Frequency of Conversion

Time

​$9010.00

5.6%

semi annually

77

years
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statement showing Computations
Particulars Amount
Principal                9,010.00
Nominal rate 5.60%
Semi Annual rate = 5.6%/2 2.80%
Time in years                      77.00
Semi Annual years = 77*2                    154.00
A= P(1+r/100)^n
Future Value =9010 *(1 + 2.80/100)^154
Future Value =9010 *(1.028)^154
Future Value =9010 *70.2975
Future Value = $633,380.15